OpenStudy (anonymous):

A rectangular prism has a length of 5 centimeters, a width of 5 centimeters, and a height of 10 centimeters. The surface area of this prism is 250 square centimeters. True or False

hero (hero):

surface area of a prism = 2(xy) + 2(xz) + 2(yz) In this case, x = 5, y = 5, and z = 10: SA = 2(5*5) + 2(5*10) + 2(5*10) SA = 50 + 100 + 100 SA = 250 True

OpenStudy (anonymous):

false